OpenStudy (anonymous):

13. W+3>9 = w>6 14. V-6<4 = v<10 15. -4

OpenStudy (saifoo.khan):

w+3>9 subtract 3 from both sides, x > 9 -3 x > 6

OpenStudy (anonymous):

Okay and what about the other 2?

OpenStudy (saifoo.khan):

add 6 on both sides.

OpenStudy (anonymous):

can you show me how to please?

OpenStudy (saifoo.khan):

v-6<4 add 6 on both sides, v < 4 + 6 v < 10

OpenStudy (anonymous):

Okay and the last one?

OpenStudy (saifoo.khan):

try that one. :)

OpenStudy (anonymous):

Add 8 on each side? -4 <t+8 Add 8 on both sides t<4+8 T=8 ?

OpenStudy (saifoo.khan):

-4 <t+8 Subtract 8 on each sides -4 -8 <t

OpenStudy (anonymous):

so when the | < | Arrow Faces that way its subtraction and the other way is addition

OpenStudy (saifoo.khan):

Nope, actually u have to bring the variable on one side of the equation. that's it.

OpenStudy (anonymous):

im sorry but im kind of confused :/ uhmm can you give and example?

OpenStudy (saifoo.khan):

x + 2 = 8 how u are going to solve it? bring the variable, the x on one side of the equation right? so, x+2 = 8 subtract 2 from both sides, x + 2 - 2 = 8 -2 x + 0 = 6 x = 6

OpenStudy (anonymous):

so each side you subtracted the 2 from can be something else? like x+2-2 and 8 -2